we expect a cars gas mileage to be related to its city mileage. data for all 1198 vehicles in the governments 2008 Fuel Economy Guide give the regression line:

highway mpg = 4.62 + (1.109 x city mpg) for predicting highway mileage from city mileage.

A.

What is the slope of this line?

Say in words what the numerical value of the slope tells us.

B. What is the intercept? Explain why the value of the intercept is not statically meaningful

C. Find the predicted highway mileage for a car that gets 16 miles per gallon in the city.

Find the predicted highway mileage for a car that gets 28 miles per gallon in the city.

PLEASE HELP EXPLAIN HOW TO DO THIS AND SHOW WORK BECAUSE I AM CONFUSED AND WANT TO KNOW THANK YOU GOD BLESS
Answer #1

A. the slope is 1.109

the increase of highway mpg 1.109 by the increase of city mpg 1.

B. the intercept is 4.62. this value is not statistically meaningful because of if city mpg is zero the highway mpg will be 4.62 + 1.109 * 0 = 4.62 is so less or not meaningful because 4.62 mpg is impossible in highway

C.highway mpg = 4.62 + (1.109 x city mpg)

highway mpg = 4.62 + 1.109 * 16

= 22.364

highway mpg = 4.62 + 1.109 * 28

= 35.672